OpenStudy (anonymous):

please help!!!when the expression 4x^2-3x-8 is divided by x-a the remainder is 2. Find the value of a

sam (.sam.):

Easy, p(x)=4x^2-3x-8 when x-a is a factor of p(x), then p(a) is equal to 2, so p(a)=4a^2-3a-8=2 Solve for 4a^2-3a-8=2

OpenStudy (anonymous):

what do you mean by solve for 4a^2-3a-8=2?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

someone help!

OpenStudy (austinl):

\(\large{4a^2-3a-8=2}\) Solve for a.

OpenStudy (anonymous):

???

sam (.sam.):

Solve for 'a', Like this 4a^2-3a-8=2 subtract 2 to both sides 4a^2-3a-10=0 Then factor it

OpenStudy (anonymous):

sorry,, my browser crashed,,

OpenStudy (anonymous):

(-5 + -4a)(2 + -1a) = 0

OpenStudy (anonymous):

next???

OpenStudy (anonymous):

(4a-5)(a+2)

OpenStudy (anonymous):

what's next??

OpenStudy (anonymous):

get the roots??

sam (.sam.):

Careful, its not (4a-5)(a+2)

OpenStudy (anonymous):

sorry,,

OpenStudy (anonymous):

xDD,,Careful, its not (4a+5)(a-2)

sam (.sam.):

Okay then (4a+5)(a-2)=0 a=-5/4 and a=2 done

OpenStudy (anonymous):

ahha.,xDD sorry,, just copy pasted.. xDD

OpenStudy (anonymous):

thanks,, so much.,